If the distance between earth and moon is 60times the radius of earth. if acceleration due to gravity on earth is 10 m/sec² find the centripetal acceleration of moon.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Accln. for earth,
g = 10 = GM/R^2
and
Accln. for moon,
g’ = Gm/(60R)^2
dividing them we get,
10/g’ = 3600M/m
g’ = m/360M m/s^2.
M is earth mass and m is moon’s mass.